Devise a synthesis of the following molecules. Rememeber to count carbons.
OO HN
NH2 NO2
Answer #1

HN Nitration of anomahic ing (Euchophilic anomatic substituhion HNO3 NO2 КMпоч, H+ axidahon of Buhzyllic carb0h No2 NO2 OH EsNO2 NO2 NO2 N N As03 -OH O NaoH NH N=N -OH OH N=N HNOS He NEBCIO -OH N N HBF -HC JEN BF ece OH N=N Oh N=N- NaNo Cu NEN BF NO